OpenStudy (anonymous):

isn't Bayes' theorem all about conditional probability and isn't actually the conditional probability?

OpenStudy (anonymous):

@satellite73 @TuringTest any help please?

OpenStudy (anonymous):

baye's theorem is just some reworking the algebra for \(P(A|B)=\frac{P(A\cap B)}{P(B)}\)at least in its most basic form

OpenStudy (anonymous):

sorry I didn't get it

OpenStudy (anonymous):

since \(P(A\cap B)=P(B|A)P(A)\) and \(P(B)=P(A\cap B) +P(A^c\cap B)=P(B|A)P(A)+P(B|A^c)P(A^c)\) you get bayes by substitution

OpenStudy (anonymous):

for two event bayes says \[P(A|B)=\frac{P(B|A)P(A)}{P(B|A)P(A)+P(B|A^c)P(A^c)}\]

OpenStudy (anonymous):

it is just an algebraic reworking of the definition, but now the conditions have switched

OpenStudy (anonymous):

so yes, it is in fact the conditional probability as you said

OpenStudy (anonymous):

@satellite73 so Baye's thm is the basics/fundamental of what we study today in the modern books of probability about conditional probability?

OpenStudy (agent47):

It's easy to understand if you use venn diagrams

OpenStudy (anonymous):

I do understand what is Baye's theoram in general (I mean how it works) I believe that is sort of the fundamental of modern conditional probability but I want to get suggestions from others so that I can see myself if I am wrong or not